FloorMate Carpet Company manufactures carpets. Fiber is placed in process in the Spinning Department, where it is spun into yarn. The output of the Spinning Department is transferred to the Tufting Department, where carpet backing is added at the beginning of the process and the process is completed. On October 1, FloorMate Carpet Company had the following inventories:

Finished Goods $5,600
Work in Process-Spinning Department 1,500
Work in Process-Tufting Department 2,300
Materials 4,800

Departmental accounts are maintained for factory overhead, and both have zero balances on October 1. Manufacturing operations for October are summarized as follows:

Oct.

1 Materials purchased on account $ 84,900
2 Materials requisitioned for use:
Fiber—Spinning Department $ 43,600
Carpet backing—Tufting Department 34,100
Indirect materials—Spinning Department 3,200
Indirect materials—Tufting Department 2,900
31 Labor used:
Direct labor—Spinning Department $ 26,300
Direct labor—Tufting Department 17,900
Indirect labor—Spinning Department 12,100
Indirect labor—Tufting Department 11,700
31 Depreciation charged on fixed assets:
Spinning Department $ 5,300
Tufting Department 3,400
31 Expired prepaid factory insurance:
Spinning Department $ 1,200
Tufting Department 1,000
31 Applied factory overhead:
Spinning Department $ 22,000
Tufting Department 18,700
31 Production costs transferred from Spinning Department to Tufting Department $ 88,000
31 Production costs transferred from Tufting Department to Finished Goods $159,000
31 Cost of goods sold during the period $160,500
Required:
1. Journalize the entries to record the operations, using the dates provided with the summary of manufacturing operations. Refer to the Chart of Accounts for exact wording of account titles.
2. Compute the October 31 balances of the inventory accounts.
3.

Compute the October 31 balances of the factory overhead accounts.

2. Compute the October 31 balances of the inventory accounts.

Materials selector 1Credit

Debit

Credit

Work in Process
-Spinning Department selector 2Credit

Debit

Credit

-Tufting Department selector 3Credit

Debit

Credit

Finished Goods selector 4Credit

Debit

Credit

Points:

Feedback

Check My Work

Explanation

3. Compute the October 31 balances of the factory overhead accounts.

Factory Overhead
-Spinning Department selector 1Debit

Debit

Credit

-Tufting Department selector 2Credit

Debit

Credit
Answer #1
Concepts and reason

Transaction: It is an exchange of value in which each participant receives or sacrifices value. It creates a legal obligation between two or more parties for exchanging the goods or providing the services. The transactions which can be estimated in money are to be noted in financial books.

Rules of debit and credit have been followed for journalizing the various transactions and these are:

Debit the Receiver and Credit the Giver: It is used for personal accounts. It indicates when the organization receives something from anyone then, what is received would be debited and the giver would be credited.

Debit what comes in and Credit what goes out: It is used for real accounts. It indicates when the organization purchased or receives any asset then it would be debited and on the other side, when the asset is going out of the organization then, it would be credited.

Debit all expenses and losses, credit all incomes and gains: It is used in case of a nominal account, and according to this rule all the expenses and losses incurred by the organization would be debited and on the other side, all the incomes and gain of the organization would be credited.

Process Costing: It is one of the types of operating costing system. This system is used in collecting and assigning the manufacturing cost to the units manufactured. This system also helps in computing the cost when the product is manufactured through a continuous process.

Fundamentals

Journal Entry: Journalizing is the process of recording of transactions in the book of original entry. It gives a complete picture of a business transaction. It is recorded in chronological order. It is the pre-phase for the preparation of ledgers.

Assets: It can be defined as the resources which are controlled and owned by the organization and which are capable of providing some future benefits for operating the core business of the organization. Examples of assets are Furniture, Machinery, Car, Cash and many others.

Liabilities: These are referred as company’s financial obligation arising during the course of a Business Cycle, which will be settled by the outflow of Resources. The money or the service owed by one party to another party is also referred as a liability.

Revenues: It refers the income that an entity earns from its normal business operations. It reflects the amount which is received by the firm by manufacturing or providing goods and services. Whenever income has been earned or due, the revenue account will be credited.

Expenses: These are the gross outflows incurred by a business entity while carrying out their regular business operations like manufacturing and providing goods and services. Whenever any expense has been due to be incurred, the expense account has been debited.

Raw Material (RM): The material of a company which is unprocessed and will be used to manufacture finished products are called raw materials. They may be direct or indirect.

Finished Goods (FG): The raw materials are processed and such goods which complete the process of manufacturing, but are not yet sold to the customer, are known as finished goods. It also includes goods in a complete form have been purchased for the purposes of sale.

Work-in-progress (WIP) Inventory: The inventory at which some percentage of production activity has been performed is WIP inventory. It is the inventory which is not 100% complete for final consumption.

Factory Overheads: It is the aggregate of Indirect Material, Indirect Labor and Indirect Expenses which are incurred during Production of Finished Goods in the Factory. These are the cost which cannot directly belong to a particular unit of a product.

Cost of goods sold: COGS are the cost used in the production of goods. It includes the cost of material, labor and overheads. Indirect expenses are excluded from COGS. It is computed by deducting the closing inventory of goods from the cost of goods which are available for sale during the year.

Actual Manufacturing Overhead Incurred: It refers as the actual cost of overhead incurred by the organization to accomplish the goal of the business by producing the finished goods. Actual manufacturing overhead has debit balance as it is an expense.

Manufacturing Overhead applied: Manufacturing overheads are applied to the production of the finished goods on the basis of the predetermined rate of overhead. Applied manufacturing overhead has credit balance because it is applied to the work-in-progress.

Total manufacturing cost: It refers the cost which is incurred to produce the products. Total manufacturing cost is the combination of all material used, factory overheads, production overheads utilized in a particular reporting period.

(1)

On Oct 1, Company purchased material on the account and the entry for the same is given below:

JOURNAL
Date
Account Titles and Explanation
1st Oct Material
Accounts payable
(To record the purchase of material)
Post Ref D

On Oct 2, the departments of company use material and the entry for the same is given below:

JOURNAL
Date Account Titles and Explanation
2nd Oct Work in progress-S Department
Work in progress- T Department
Factory Over

On October 31, Company charged labor from different departments and the entry for the same is given below:

JOURNAL
Date | Account Titles and Explanation
31st Oct Work in progress- S Department
Work in progress- T Department
Factory

On October 31, Company record the depreciation expense of each department and the entry for the same is given below:

JOURNAL
Date Account Titles and Explanation
31st Oct Factory Overhead-S Department
Factory Overhead-T Department
Accumulated

On October 31, Company records the factory insurance expense and the entry for the same is given below:

JOURNAL
Date Account Titles and Explanation
31st Oct Factory Overhead-S Department
Factory Overhead-T Department
Prepaid Insu

On October 31, Company record the factory overhead applied and the entry for the same is given below:

JOURNAL
| Date
Account Titles and Explanation
31st Oct Work in progress-S Department
Work in progress- T Department
Factory O

On October 31, Company records the transfer of production cost and the entry for the same is given below:

JOURNAL
Date
Account Titles and Explanation
31st Oct Work in progress- T Department
Work in progress- S Department
(To record

On October 31, Company transfer the production cost to the finished good and the entry for the same is given below:

JOURNAL
Date Account Titles and Explanation
| 31st OctFinished goods Inventory
Work in progress- T Department
|(To record the

On October 31, Company records the cost of goods sold and the entry for the same is given below:

JOURNAL
Date
Account Titles and Explanation
31st Oct Cost of goods sold
Finished goods Inventory
(To record the cost of goods

(2)

Compute the ending balance of material:

Particulars
Balance as on October 1
Add: Purchases
Total cost of materials available
Less: Materials requisitioned
Balance as

Working notes:

Compute the total cost of material available:

Totalcostofmaterialavailable=(BalanceasonOctober1+PurchasesduringOctober)=($4,800+$84,900)=$89,700\begin{array}{c}\\{\rm{Total}}\,{\rm{cost}}\,{\rm{of}}\,{\rm{material}}\,{\rm{available}} = \left( {{\rm{Balance}}\,{\rm{as}}\,{\rm{on}}\,{\rm{October}}\,1 + {\rm{Purchases}}\,{\rm{during}}\,{\rm{October}}} \right)\\\\ = \left( {\$ 4,800 + \$ 84,900} \right)\\\\ = \$ 89,700\\\end{array}

Therefore, the total cost of material available is $89,700.

Compute the balance of material as on October 31:

BalanceofmaterialOctober31=(TotalcostofmaterialavailableMaterialrequisitioned)=($89,700$83,800)=$5,900\begin{array}{c}\\{\rm{Balance}}\,{\rm{of}}\,{\rm{materia}}{{\rm{l}}_{{\rm{October}}\,31}} = \left( {{\rm{Total}}\,{\rm{cost}}\,{\rm{of}}\,{\rm{material}}\,{\rm{available}} - Material\,requisitioned} \right)\\\\ = \left( {\$ 89,700 - \$ 83,800} \right)\\\\ = \$ 5,900\\\end{array}

Therefore, the ending balance of material is $5,900.

Calculate the balance of work-in-progress of S Department as on October 31:

Amount ($)
1,500
Particulars
Balance as on October 1
Add: Manufacturing cost added during
October:
Materials
Direct Labor
Fac

Working notes:

Compute the total manufacturing cost:

Totalmanufacturingcost=(BalanceOctober1+Material+Directlabor+Factoryoverhead)=($1,500+$43,600+$26,300+$22,000)=$93,400\begin{array}{c}\\{\rm{Total}}\,{\rm{manufacturing}}\,{\rm{cost}} = \left( \begin{array}{l}\\{\rm{Balanc}}{{\rm{e}}_{{\rm{October}}\,1}} + {\rm{Material}} + {\rm{Direct}}\,{\rm{labor}}\\\\ + {\rm{Factory}}\,{\rm{overhead}}\\\end{array} \right)\\\\ = \left( {\$ 1,500 + \$ 43,600 + \$ 26,300 + \$ 22,000} \right)\\\\ = \$ 93,400\\\end{array}

Therefore, the total manufacturing cost is $93,400.

Compute the balance of work-in-progress as on October 31:

BalanceofWIPOct31=(TotalmanufacturingcostManufacturingcosttransfertoTdepartment)=($93,400$88,000)=$5,400\begin{array}{c}\\{\rm{Balance}}\,{\rm{of}}\,{\rm{WI}}{{\rm{P}}_{{\rm{Oct}}\,31}} = \left( \begin{array}{l}\\{\rm{Total}}\,{\rm{manufacturing}}\,{\rm{cost}}\\\\ - {\rm{Manufacturing}}\,{\rm{cost}}\,{\rm{transfer}}\,{\rm{to}}\,{\rm{T}}\,{\rm{department}}\\\end{array} \right)\\\\ = \left( {\$ 93,400 - \$ 88,000} \right)\\\\ = \$ 5,400\\\end{array}

Therefore, the balance of WIP as on October 31 is $5,400.

Calculate the balance of work-in-progress of T Department as on October 31:

Amount (3)
2,3001
Particulars
Balance as on October 1
Add: Manufacturing cost added during
October:
Materials
Direct Labor
Fa

Working notes:

Compute the total manufacturing cost:

Totalmanufacturingcost=(BalanceOctober1+Material+Directlabor+Factoryoverhead+ManufacturingcostfromSDepartment)=($2,300+$34,100+$17,900+$18,700+$88,000)=$161,000\begin{array}{c}\\{\rm{Total}}\,{\rm{manufacturing}}\,{\rm{cost}} = \left( \begin{array}{l}\\{\rm{Balanc}}{{\rm{e}}_{{\rm{October}}\,1}} + {\rm{Material}} + {\rm{Direct}}\,{\rm{labor}} + {\rm{Factory}}\\\\{\rm{overhead + Manufacturing}}\,{\rm{cost}}\,{\rm{from S Department}}\\\end{array} \right)\\\\ = \left( {\$ 2,300 + \$ 34,100 + \$ 17,900 + \$ 18,700 + \$ 88,000} \right)\\\\ = \$ 161,000\\\end{array}

Therefore, the total manufacturing cost is $161,000.

Compute the balance of work-in-progress as on October 31:

BalanceofWIPOct31=(TotalmanufacturingcostManufacturingcosttransfertoFinishedgoods)=($161,000$159,000)=$2,000\begin{array}{c}\\{\rm{Balance}}\,{\rm{of}}\,{\rm{WI}}{{\rm{P}}_{{\rm{Oct}}\,31}} = \left( \begin{array}{l}\\{\rm{Total}}\,{\rm{manufacturing}}\,{\rm{cost}}\\\\ - {\rm{Manufacturing}}\,{\rm{cost}}\,{\rm{transfer}}\,{\rm{to}}\,{\rm{Finished goods}}\\\end{array} \right)\\\\ = \left( {\$ 161,000 - \$ 159,000} \right)\\\\ = \$ 2,000\\\end{array}

Therefore, the balance of WIP as on October 31 is $2,000.

Compute the balance of finished goods as on October 31:

Amount (S)
5,600
Particulars
Balance as on October 1
Add: Manufacturing cost transfer from
T Department
Total cost of finishe

Working notes:

Compute the total cost of finished goods available for sale:

Totalcostoffinishedgoodsavailableforsale=(BalanceasonOctober1+ManufacturingcostfromTdepartment)=($5,600+$159,000)=$164,600\begin{array}{c}\\{\rm{Total}}\,{\rm{cost}}\,{\rm{of}}\,{\rm{finished}}\,{\rm{good}}{{\rm{s}}_{{\rm{available}}\,{\rm{for}}\,{\rm{sale}}}} = \left( \begin{array}{l}\\{\rm{Balance}}\,{\rm{as}}\,{\rm{on}}\,{\rm{October}}\,1\\\\ + {\rm{Manufacturing}}\,{\rm{cost}}\,{\rm{from}}\,{\rm{T}}\,{\rm{department}}\\\end{array} \right)\\\\ = \left( {\$ 5,600 + \$ 159,000} \right)\\\\ = \$ 164,600\\\end{array}

Therefore, the total cost of finished goods available for sale is $4,100.

Compute the balance of finished goods as on October 31:

BalanceofFinishedgoodsOctober31=(TotalcostoffinishedgoodsavailableforsaleCostofgoodssold)=($164,600$160,500)=$4,100\begin{array}{c}\\{\rm{Balance}}\,{\rm{of}}\,{\rm{Finished}}\,{\rm{good}}{{\rm{s}}_{{\rm{October}}\,31}} = \left( \begin{array}{l}\\{\rm{Total}}\,{\rm{cost}}\,{\rm{of}}\,{\rm{finished}}\,{\rm{goods}}\,{\rm{available}}\,{\rm{for}}\,{\rm{sale}}\\\\ - {\rm{Cost}}\,{\rm{of}}\,{\rm{goods}}\,{\rm{sold}}\\\end{array} \right)\\\\ = \left( {\$ 164,600 - \$ 160,500} \right)\\\\ = \$ 4,100\\\end{array}

Therefore, the balance of finished goods as on October 31 is $4,100.

(3)

Calculate the balance of Factory overhead-S Department as on October 31:

Amount (S)
Particulars
Balance as on October 1
Indirect materials
Indirect labor
Depreciation charged on fixed assets
Expired

Working notes:

Compute total factory overhead cost incurred:

TotalFactoryoverheadcostincurred=(Indirectmaterial+Directlabor+Depreciationexpense+FactoryInsurance)=($3,200+$12,100+$5,300+$1,200)=$21,800\begin{array}{c}\\{\rm{Total}}\,{\rm{Factory}}\,{\rm{overhead}}\,{\rm{cost incurred}} = \left( \begin{array}{l}\\{\rm{Indirect}}\,{\rm{material}} + {\rm{Direct}}\,{\rm{labor}}\\\\ + {\rm{Depreciation}}\,{\rm{expense}} + {\rm{Factory}}\,{\rm{Insurance}}\\\end{array} \right)\\\\ = \left( {\$ 3,200 + \$ 12,100 + \$ 5,300 + \$ 1,200} \right)\\\\ = \$ 21,800\\\end{array}

Therefore, the total factory overhead cost incurred is $21,800.

Compute the balance of factory overhead-S department as on October 31:

FactoryOverheadOctober31=(TotalfactoryoverheadcostincurredFactoryoverheadapplied)=($21,800$22,000)=$200(Creditbalance)\begin{array}{c}\\{\rm{Factory}}\,{\rm{Overhea}}{{\rm{d}}_{{\rm{October}}\,31}} = \left( \begin{array}{l}\\{\rm{Total}}\,{\rm{factory}}\,{\rm{overhead}}\,{\rm{cost}}\,{\rm{incurred}}\\\\ - {\rm{Factory}}\,{\rm{overhead}}\,{\rm{applied}}\\\end{array} \right)\\\\ = \left( {\$ 21,800 - \$ 22,000} \right)\\\\ = - \$ 200\,\left( {{\rm{Credit}}\,{\rm{balance}}} \right)\\\end{array}

Therefore, the balance of factory overhead as on October 31 is -$200.

Calculate the balance of Factory overhead-T Department as on October 31:

Amount (S)
Particulars
Balance as on October 1
Indirect materials
Indirect labor
Depreciation charged on fixed assets
Expired

Working notes:

Compute total factory overhead cost incurred:

TotalFactoryoverheadcostincurred=(Indirectmaterial+Directlabor+Depreciationexpense+FactoryInsurance)=($2,900+$11,700+$3,400+$1,000)=$19,000\begin{array}{c}\\{\rm{Total}}\,{\rm{Factory}}\,{\rm{overhead}}\,{\rm{cost incurred}} = \left( \begin{array}{l}\\{\rm{Indirect}}\,{\rm{material}} + {\rm{Direct}}\,{\rm{labor}}\\\\ + {\rm{Depreciation}}\,{\rm{expense}} + {\rm{Factory}}\,{\rm{Insurance}}\\\end{array} \right)\\\\ = \left( {\$ 2,900 + \$ 11,700 + \$ 3,400 + \$ 1,000} \right)\\\\ = \$ 19,000\\\end{array}

Therefore, the total factory overhead cost incurred is $19,000.

Compute the balance of factory overhead-T department as on October 31:

FactoryOverheadOctober31=(TotalfactoryoverheadcostincurredFactoryoverheadapplied)=($19,000$18,700)=$300(Debitbalance)\begin{array}{c}\\{\rm{Factory}}\,{\rm{Overhea}}{{\rm{d}}_{{\rm{October}}\,31}} = \left( \begin{array}{l}\\{\rm{Total}}\,{\rm{factory}}\,{\rm{overhead}}\,{\rm{cost}}\,{\rm{incurred}}\\\\ - {\rm{Factory}}\,{\rm{overhead}}\,{\rm{applied}}\\\end{array} \right)\\\\ = \left( {\$ 19,000 - \$ 18,700} \right)\\\\ = \$ 300\,\left( {{\rm{Debit}}\,{\rm{balance}}} \right)\\\end{array}

Therefore, the balance of factory overhead as on October 31 is $300.

Ans: Part 1.1

JOURNAL
Date
Account Titles and Explanation
1st Oct Material
Accounts payable
(To record the purchase of material)
Post Ref D

Part 1.2

JOURNAL
Date Account Titles and Explanation
2nd Oct Work in progress-S Department
Work in progress- T Department
Factory Over

Part 1.3

JOURNAL
Date | Account Titles and Explanation
31st Oct Work in progress- S Department
Work in progress- T Department
Factory

Part 1.4

JOURNAL
Date Account Titles and Explanation
31st Oct Factory Overhead-S Department
Factory Overhead-T Department
Accumulated

Part 1.5

JOURNAL
Date Account Titles and Explanation
31st Oct Factory Overhead-S Department
Factory Overhead-T Department
Prepaid Insu

Part 1.6

JOURNAL
| Date
Account Titles and Explanation
31st Oct Work in progress-S Department
Work in progress- T Department
Factory O

Part 1.7

JOURNAL
Date
Account Titles and Explanation
31st Oct Work in progress- T Department
Work in progress- S Department
(To record

Part 1.8

JOURNAL
Date Account Titles and Explanation
| 31st OctFinished goods Inventory
Work in progress- T Department
|(To record the

Part 1.9

JOURNAL
Date
Account Titles and Explanation
31st Oct Cost of goods sold
Finished goods Inventory
(To record the cost of goods

Part 2.1

The balance of material as on October 31 is $5,900.

Part 2.2

The balance of WIP-S Department as on October 31 is $5,400.

Part 2.3

The balance of WIP-T Department as on October 31 is $2,000.

Part 2.4

The balance of finished goods inventory as on October 31 is $4,100.

Part 3.1

The balance of factory overhead-S department as on October 31 is -$200.

Part 3.2

The balance of factory overhead-T department as on October 31 is $300.